OpenStudy (ajprincess):

Pls help:) Solve each of the following: \(a)4y''+4y'+37y=5xe^{-2x}\) \(b)y"+4y=12x^2-16xcos2x\)

OpenStudy (ajprincess):

For the a) part I found the complimentary function this way. Given \((4D^2+4D+37)y=5xe^{-2x}\) In auxiliary equation is \(4\lambda ^2+4\lambda +37=0\) \[\lambda=\frac{-1}{2}\pm3i\] Thus, the complimentary function is \[y_n=e^{\frac{-1x}{2}}(C_1cos3x+C-2sin3x)\]

OpenStudy (ajprincess):

@wio hav I done it correctly?.

OpenStudy (anonymous):

I'm not sure actually.

OpenStudy (ajprincess):

oh k. Are there any other methods for doing this

OpenStudy (anonymous):

It's been a while sense I've done 2nd Order with complex roots.

OpenStudy (anonymous):

By complementary solution, you mean the homogeneous solution where we set it equal to 0?

OpenStudy (ajprincess):

ya.:)

OpenStudy (anonymous):

Well, the \(C-2\) threw me off for a while, but realizing it is \(C_2\), I'd say you are correct.

OpenStudy (anonymous):

Next is the particular solution.

OpenStudy (ajprincess):

ya. that's where I am stuck.

OpenStudy (anonymous):

Yeah, actually normally I would say guess and check, but I've recently learned there may be a more elegant method.

OpenStudy (anonymous):

http://math.mit.edu/suppnotes/suppnotes03/o.pdf

OpenStudy (ajprincess):

hmm. i did a similar question with real roots. but am finding it hard to do it with the complex one.

OpenStudy (anonymous):

Hmmm, how did it go with real roots?

OpenStudy (anonymous):

I really don't see a difference, since you just add the homogeneous solution to the particular one.

OpenStudy (ajprincess):

\(y""-3y'+2y=xe^x\) this is the question. The particular solution was found this way. \[Y_p=\frac{1}{D^2-3D+2}xe^x\] \[=\frac{1}{(D-1)(D-2)}xe^x\] \[=\frac{1}{(D-1)}e^{2x}\int e^{-2x}xe^xdx\] \[=\frac{1}{(D-1)}e^{2x}[-e^{-x}-\int e^{-x}dx\] \[=\frac{1}{(D-1)}[-(x+1)e^x]\] \[=-e^x\int e^x(x+1)e^xdx\] \[=\frac{-1}{2}(1+x)^2e^x\]

OpenStudy (anonymous):

So you didn't use undetermined coefficients... interesting. Yet this method would be tricky for complex roots since it can't be properly factored.

OpenStudy (anonymous):

The undetermined coefficients method is to try: \[ Ce^{-2x}(Ax+B) \]and then solve for the coefficients.

OpenStudy (anonymous):

To be honest, I never learned your method. But I think it could be extended with Euler's formula. If I understood it.

OpenStudy (anonymous):

Looks like the pattern is: \[ \frac{1}{D-a}f(x) = e^{ax}\int e^{-ax}f(x)dx \]

OpenStudy (ajprincess):

ohh. Actually I missed this lecture. somehw i understood this nd can do for questions with real roots. but am unable to do for the complex ones. ya that's the general formula given to us.

OpenStudy (anonymous):

I noticed that on the final step with \(D-1\) the inner \(e^x\) doesn't have the negative, but according to the pattern it should...?

OpenStudy (anonymous):

@ajprincess Can you try just doing it with complex number numbers?

OpenStudy (anonymous):

Just let \(a\) be complex. Let's see the final result I guess.

OpenStudy (ajprincess):

Really sorry.probs in the internet connection. in the same way as I did for real roots.

OpenStudy (ajprincess):

@wio do I hav to do it in the same way?

OpenStudy (anonymous):

If I were you I would look up the pattern in the book I would look it up myself if I knew the book and where you were.

OpenStudy (anonymous):

And if you find the pattern in the book, post it here so we can learn.

OpenStudy (ajprincess):

Ha k. sure:)

OpenStudy (anonymous):

I found something interesting that may help.

OpenStudy (anonymous):

\[ \frac{1}{f(D)}e^{ax}f(x) = e^{ax}\frac{1}{D+a}f(x) \]

OpenStudy (anonymous):

If we shifted it over, it might end up having real roots

OpenStudy (anonymous):

http://www.solitaryroad.com/c657.html

OpenStudy (ajprincess):

 do u mean this @wio?